# Principal stratification with $U$-statistics under principal ignorability

### v1

@@ 5 Simulation studies
We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 6, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15). In each simulation setting, we consider two scenarios for each of the three nuisance functions, i.e., whether it is correctly specified, leading to a total of eight different combinations.
@@ 5.1 The probability index estimands with an intermediate variable
For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 4. For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner package in R with the random forest and generalized linear models libraries. We set the number of observations $n=2,000$ and consider $1,000$ replications; violin plots containing the simulation results under each considered estimator are given in Figure 1 and Web Figure 1 in Section S6 of the Online Supplement.
@@ 5.2 The win-ratio estimands with an intermediate variable
In the second simulation study, we consider the case of the win-ratio estimands introduced in Examples 2 and 5, with the final potential outcomes being ordinal. We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study. However, for the potential outcome $Y_{i}(z)\in\{1,\ldots,Q\}$, we consider the following three-category ordinal logistic regression model:
@@ 6.1 The survivor probability index for continuous outcomes
The estimated results for SPI are given in Table 1, where $1,000$ bootstrap replicates were obtained for estimating the standard errors. We observe that, for the always-employed (72.9% of all participants by estimation), the probability index is estimated as $0.539$ using the triply robust and the debiased machine learning methods. This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-survivor under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment. Finally, the point estimates under the plug-in moment methods are similar to the triply robust and debiased machine learning methods, which likely suggests that each parametric working model may be approximately correctly specified.

### v2

@@ 5 Simulation studies
We conducted simulation studies to evaluate the finite-sample performance of the proposed estimators for PGCEs, including the three plug-in moment estimators motivated by Theorem 3.4, the triply robust estimators with parametric working models given in (14), and the debiased machine learning estimators given in (15). In each simulation setting, we consider two scenarios for each of the three nuisance functions, i.e., whether it is correctly specified, leading to a total of eight different combinations.
@@ 5.1 The probability index estimands with an intermediate variable
For the plug-in moment estimators and the triply robust estimator with parametric working models, treatment propensity scores and principal scores are estimated via logistic regression, while the pairwise outcome mean is estimated via linear regression as in Example 3.2. For the debiased machine learning estimator, we adopt a five-fold cross-fitting ($K=5$ and $L=15$) and estimate nuisance functions using the SuperLearner package in R with the random forest and generalized linear models libraries. We set the number of observations $n=2,000$ and consider $1,000$ replications; violin plots containing the simulation results under each considered estimator are given in Figure 1 and Web Figure 1 in Section S6 of the Online Supplement.
@@ 5.2 The win-ratio estimands with an intermediate variable
In the second simulation study, we consider the case of the win-ratio estimands introduced in Examples 2.2 and 3.3, with the final potential outcomes being ordinal. We generate covariates $X_{i}$, the treatment assignment $Z_{i}$, and the intermediate variable $D_{i}$ from the same process given in the previous simulation study. However, for the potential outcome $Y_{i}(z)\in\{1,\ldots,Q\}$, we consider the following three-category ordinal logistic regression model:
@@ 6.1 The survivor probability index for continuous outcomes
The estimated results for SPI are given in Table 1, where $1,000$ bootstrap replicates were obtained for estimating the standard errors. We observe that, for the always-employed (72.9% of all participants by estimation), the probability index is estimated as $0.539$ using the triply robust and the debiased machine learning methods. This indicates that the probability that a randomly chosen always-employed under assigned treatment has a higher-ranked earning outcome than a randomly chosen always-employed under control is $0.539$, with the 95% confidence interval excluding the null of $0.5$; this implies that the Jobs Corps positively affects one’s earnings in the one year after the assignment. Finally, the point estimates under the plug-in moment methods are similar to the triply robust and debiased machine learning methods, which likely suggests that each parametric working model may be approximately correctly specified.